If ‘s’ is the saving ratio of a country, and ‘k’ is the capital-output ratio, then the GDP growth rate is represented by ‘k/s’. True or False. Explain your answer briefly and precisely.


Assume that a country’s GDP must grow by 2% per year. Now, the country’s capital-output ratio (k) is 6. To achieve the required growth rate 2%, what must the savings ratio of this country? Explain briefly.
Answer #1

Capital Output Ratio explains the relationship between investments and Economic Growth of a Country. In other words, it is the capital needed to produce one unit of Output. So If K is the Capital Output Ratio and s is the saving ratio and let's assume g is the growth

Therefore, G = s/k and hence G = k/s is false.

Here.

G = 2%

K = 6

To remain at 2%, saving Ratio required is, 2% = s/ 6

There s = 12%